OpenStudy (mayankdevnani):

My first TUTORIAL on Slope Intercept Form """""::::

hartnn (hartnn):

*waiting for it*

OpenStudy (mayankdevnani):

\[\Huge{\bf{\color{blue}{Slope Intercept }\color{red}{Form--}}}\] \[\Large{\bf{\color{blue}{SlopeIntercept}\color{red}{Form-}}}\] is also called as gradient, y-intercept form. Slope Intercept Form is used to generate the Equation of a straight line with a slope m and y - intercept c of the line. \[\Huge{\bf{\color{green}{Form}\color{}{ula-}}}\] Slope Intercept Form : Slope Intercept Form = y = mx + c where c = y-intercept of the line, m = Slope of the line \[\Huge{\bf{\color{blue}{SlopeIntercept}\color{grey}{Diagram--}}}\]  \[\Huge{\bf{\color{red}{SlopeIntercept}\color{green}{Example--}}}\] Determine the equation of the line with slope of 3 and y-intercept 2. slope m = 3. slope c = 2. Substitute in the formula as : y = mx + c y = (3)x + (2) y = 3x + 2 Equation = 3x + 2 -y Equation of the Line = 3x -y + 2 The above example will clearly illustrates how to calculate the Slope Intercept Form manually.

OpenStudy (mayankdevnani):

@hartnn @Hero @hba @hashsam1 @aceace @aroub @ashna @sara1234 @Shadowys @SFAdrienne @stephielee have a look!!!!

OpenStudy (hba):

Are You done with it ?

OpenStudy (mayankdevnani):

thnx... @hartnn

OpenStudy (hba):

Anyways Good Work :)

OpenStudy (anonymous):

No offense, but see this: http://easycalculation.com/analytical/learn-slope-intercept.php

OpenStudy (mayankdevnani):

yaaa @hba

hartnn (hartnn):

O.o

OpenStudy (anonymous):

What a coincidence! The same example?!

OpenStudy (mayankdevnani):

.php is my website

hartnn (hartnn):

thats not co-incidence php cannot be your website....

OpenStudy (mayankdevnani):

why.html like .php

OpenStudy (hba):

lol :P

OpenStudy (anonymous):

haha :D

OpenStudy (hba):

@mayankdevnani ?

OpenStudy (hba):

Well Creative Commons License :D